Find the length of a ladder leaning against a building if the top of the ladder touches the building at a height of 12 feet. Also, the length of the ladder is 4 feet more than its distance from the base of the building.
Answer by checkley77(12844)

12^2+x^2=(x+4)^2
144+x^2=x^2+8x+16
x^2-x^2-8x=16-144
-8x=-128
x=-128/-8
x=16 ft. answer for the length of the base of the ladder from the building.
16+4=20 ft. is the length of the ladder.
Proof:
12^2+16^2=20^2
144+256=400
400=400